OpenStudy (tennistar):

What are the solutions of 2x2 + 8x = –26?

OpenStudy (thecuteone25):

do you know how to get X by it's self

OpenStudy (tennistar):

You isolate it using division or subtraction, right?

OpenStudy (thecuteone25):

right so do that and tell me what you get

OpenStudy (thecuteone25):

do division

OpenStudy (tennistar):

And divide both sides by 2, or 8. Which x are we trying to isolate?

OpenStudy (anonymous):

I can give you a hint your outcome will be a fraction.

OpenStudy (tennistar):

After dividing both sides? Or the final answer. Cause if you're saying its the final answer, you're wrong.

OpenStudy (tennistar):

These are my answer choices: –2 ± 3i –2 ± i 2 ± 6i –3 ± 2i

OpenStudy (anonymous):

my bad i was doing it a different way. Now let me try to find this way.

OpenStudy (tennistar):

No worries. Could you help me work through it as well once you solve it? I learned this earlier but I kind of forgot and need some touching up on it.

OpenStudy (anonymous):

lets do it one step at a time. 1. subtract (-26) from both sides. What would you get?

OpenStudy (anonymous):

Also there is two way on solving this. Would you like to solve it by completing the square or solving it by using the quadratic formula.

OpenStudy (tennistar):

2x2 + 8x + 26 = 0

OpenStudy (anonymous):

Ill show you both ways you can do this. -Completing the square: 1. divide both sides by 2 x^2+4x+13=0 2. subtract 13 x^2+4x=-13 3.add (4/2)^2 to both sides x^2+4x+4=-9 4. Complete the left side by using the formula x^2+2ax+a^2=(x+a)^2 (x+2)^2=-9 5. for 2 f(x)=f(a) the solution are f(x)= sq.f(a),- sq. f(a) x+2= sq. root of -9 then refine that to x+2=3i subtract 2 from both sides your answer would be x=-2+3i x= -2-3i if you don't understand that i can show you using the quadratic formula too.

OpenStudy (anonymous):

Here is the quadratic formula.